# Similarity-Distance-Magnitude Activations

### v1

@@ 6 Conclusion
We introduced $\operatorname{\text{{sdm}}}$ activation functions, which provide more robust estimators of the predictive uncertainty than the commonly-used $\operatorname{softmax}$ function. In this way, $\operatorname{\text{{sdm}}}$ activations provide a principled, data-driven substrate for approaching selective classification, calibration, and out-of-distribution detection with LLMs.
@@ Appendix A Appendix
We provide the results for the far out-of-distribution (OOD) shuffled datasets in Section A.1. Additional implementation details are included in Section A.2.
@@ A.1 Far OOD Shuffled Datasets
A.2 Additional Implementation Details
The code, data, and models to replicate and reproduce our results are available at the following URL: https://github.com/ReexpressAI/sdm_activations
We mean center the input to $g$, the 1-D CNN of the $\operatorname{\text{{sdm}}}$ activation layer and the otherwise identical CNN adaptors of the baseline comparison estimators, via the mean and standard deviation over $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$. In all experiments with adaptor layers, $M=1000$ and we use a mini-batch size of 50. We use the Adam optimizer (Kingma & Ba, 2017) with a learning rate of $1\times 10^{-5}$ for training.
A.2.1 Implementation of the SDM Activation Function
As is typical with implementations of the $\operatorname{softmax}$ function, for numerical stability, rather than directly calculating $\operatorname{\text{{sdm}}}({\bm{z}}^{\prime})$, we instead use the mathematically equivalent $\operatorname{\text{{sdm}}}({\bm{z}}^{\prime}-\max({\bm{z}}^{\prime}))$, shifting the input vector by its maximum value.
A.2.2 Implementation of the Empirical CDF Function
The conventions for implementing the empirical CDF functions follow in the expected ways, but we briefly highlight the key considerations below, as they can impact the behavior of the estimators.
  1. 1.
    The distance quantiles should be exclusionary at the boundaries. When $\operatorname{\text{$d_{\rm{nearest}}$}}=0$, the $1-{\rm{eCDF}}^{\cdot}_{\rm{ca}}(\operatorname{\text{$d_{\rm{nearest}}$}})$ quantile should be 1, and when $\operatorname{\text{$d_{\rm{nearest}}$}}$ is greater than the maximum observed distance (across $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ for ${\bm{x}}\in\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$ and ${\bm{x}}\in\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$, and across $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ for ${\bm{x}}\in\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$, the latter case only occurring during training), the $1-{\rm{eCDF}}^{\cdot}_{\rm{ca}}(\operatorname{\text{$d_{\rm{nearest}}$}})$ quantile should be 0.
  2. 2.
    For the quantiles over an $\operatorname{\text{{sdm}}}$ activation, as needed for Alg. 1, saturated values at the high-end should be assigned a quantile of 1.

### v2

@@ 6 Conclusion
We introduced $\operatorname{\text{{sdm}}}$ activation functions and $\operatorname{\text{{sdm}}}$ estimators, which are more robust estimators of the predictive uncertainty than those based on the commonly used $\operatorname{softmax}$ function. In this way, $\operatorname{\text{{sdm}}}$ activations provide a principled, data-driven substrate for approaching selective classification, calibration, and out-of-distribution detection with language models.
@@ Appendix A Appendix
We provide the results for the far out-of-distribution (OOD) shuffled datasets in Appendix A.1. Additional implementation details are included in Appendix A.2. Appendix A.3 provides an approach for analyzing the effective sample size for both the class-conditional and prediction-conditional estimates.
@@ A.1 Far OOD Shuffled Datasets
A.2 Additional Implementation Details
Replication code is available at the following URL: https://github.com/ReexpressAI/sdm_activations
We mean center the input to $g$, the 1-D CNN of the $\operatorname{\text{{sdm}}}$ activation layer and the otherwise identical CNN adaptors of the baseline comparison estimators, via the mean and standard deviation over $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$. In all experiments with adaptor layers, $M=1000$ and we use a mini-batch size of 50. We use the Adam optimizer (Kingma & Ba, 2017) (without weight decay) with a learning rate of $1\times 10^{-5}$ for training.
A.2.1 Implementation of the SDM Activation Function
As is typical with implementations of the $\operatorname{softmax}$ function, for numerical stability, rather than directly calculating $\operatorname{\text{{sdm}}}({\bm{z}}^{\prime})$, we instead use the equivalent $\operatorname{\text{{sdm}}}({\bm{z}}^{\prime}-\max({\bm{z}}^{\prime}))$, shifting the input vector by its maximum value.
A.2.2 Implementation of the Empirical CDF Function
The empirical CDF functions are assumed to be implemented such that the distance quantiles are exclusionary at the boundaries. When $\operatorname{\text{$d_{\rm{nearest}}$}}=0$, the $1-{\rm{eCDF}}^{\cdot}_{\rm{ca}}(\operatorname{\text{$d_{\rm{nearest}}$}})$ quantile should be 1, and when $\operatorname{\text{$d_{\rm{nearest}}$}}$ is greater than the maximum observed distance (across $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ for ${\bm{x}}\in\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$ and ${\bm{x}}\in\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$, and across $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ for ${\bm{x}}\in\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$, the latter case only occurring during training), the $1-{\rm{eCDF}}^{\cdot}_{\rm{ca}}(\operatorname{\text{$d_{\rm{nearest}}$}})$ quantile should be 0.
A.3 Analyzing the Effective Sample Size
In the context of the $\operatorname{\text{{sdm}}}$ estimator, to parameterize the prior belief that data points with a looser connection to $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ reflect smaller effective sample sizes, while also explicitly accounting for the count of observed points in $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$, the effective sample size for each test instance can be estimated with the following conservative assumption:
Assumption A.1.
The effective sample size is increasing in $q^{\prime}$, class-wise over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$.
For each ${\bm{x}}\in\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$, using $q^{\prime}$, we calculate $\mathbf{\hat{n}}$, the vector of effective sample sizes across classes, relative to $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$, as:
$\mathbf{\hat{n}}=\left[|\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}|^{y_{1}}\cdot{\rm{eCDF}}^{y_{1}}_{\rm{ca}}(q^{\prime}),\ldots,|\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}|^{y_{C}}\cdot{\rm{eCDF}}^{y_{C}}_{\rm{ca}}(q^{\prime})\right]$ (10)
where $|\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}|^{y_{c}}$ is the count of calibration set points with true label $y=c$.
The estimate of the effective sample size for each label can then be used to estimate the Beta-distributed error term of split-conformal coverage (Vovk, 2012), providing a sample-size-based error estimate for the class-conditional estimate, assuming exchangeability.
For the prediction-conditional estimates, assuming independent and identically distributed (i.i.d.) data, these sample size estimates can be used to construct a band around the empirical CDFs over $\operatorname{\text{$d_{\rm{nearest}}$}}$ (Eq. 5) using the sharp constant (Massart, 1990) of the distribution-free DKW inequality (Dvoretzky et al., 1956), with $\hat{n}_{\rm{min}}$ taken as the minimum among the estimated sample sizes across classes for the test instance:
$\epsilon$ $=\sqrt{\frac{1}{2\cdot\hat{n}_{\rm{min}}}\log_{e}\left(\frac{2}{1-\alpha}\right)},$ (11)
$\hat{n}_{\rm{min}}$ $=\min\left[\hat{n}_{1},\ldots,\hat{n}_{C}\right]$ (12)
If $\hat{n}_{\rm{min}}=0$, our convention is to set $\epsilon=1$. We then construct the conservative lower and upper counterparts to the distance quantile of Eq. 5:
$\operatorname{\text{$d$}}_{\rm{lower}}=\max\left(\operatorname{\text{$d$}}-\epsilon,0\right)$ (13)
$\operatorname{\text{$d$}}_{\rm{upper}}=\min\left(\operatorname{\text{$d$}}+\epsilon,1\right)$ (14)
Eq. 6 can then be calculated by substituting $\operatorname{\text{$d$}}_{\rm{lower}}$ and $\operatorname{\text{$d$}}_{\rm{upper}}$, in turn, for $\operatorname{\text{$d$}}$, resulting in a band around the prediction-conditional estimate.